SpaceRef

SpaceRef


NOAA GOES SATOPS Morning Report - August 12, 2003  Day 224 

Status Report From: National Oceanic and Atmospheric Administration
Posted: Tuesday, August 12, 2003

DMSP 

All DMSP operations were nominal over the past 24 hours. 

POES 

NOAA-15 rev 27262 / F at 2142Z on August 11:  GAC 3B data and not recorded to ground lost due to ground system problems. 

All other POES operations were nominal over the past 24 hours. 

GOES 

GOES battery reconditioning activities for all five spacecraft in process this week as part of normal pre-eclipse season preparations. 

GOES-8 continues its westward drift at 1.06 deg/day.  The drift table with the estimated current position can be found at  http://www.oso.noaa.gov/goes/GOES-8_Drift_Table.txt.  GOES-8 will be stopped at 165E on or about August 24. 

All GOES operations were nominal over the past 24 hours.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.